### v1

# Information entropy for central 197Au+197Au collisions in the UrQMD model

Abstract
Multiplicity information entropy in central 197Au + 197Au collisions at impact parameters of 03 fm are calculated at various center of mass energies () of 5.0, 7.7, 11.5, 14.5, 19.6, 27.0, 32.0, 35.0, 39.0, and 54.4 GeV using the Ultra-relativistic Quantum Molecular Dynamics model (UrQMD). The simulations in UrQMD model are compared with hydro modes with three different equations of state (EoS) and also with a default mode without hydrodynamics. The study reveals that the information entropies of baryons, net-baryons and net-protons with different equations of state in the hydro modes exhibit first decreases and then slowly increases with the increase of , while those of hadrons and anti-hadrons, antibaryons show a monotonous increase with . An enhancement is found around 30 GeV potentially corresponding to the critical endpoint with chiral hadron gas EoS and Bag model EoS.
pacs:
25.70.-z, 24.10.Lx, 21.30.FeI Introduction
The investigation of the phase diagram of quantum chromodynamics (QCD), depicting the relationship between temperature () and baryon chemical potential (), has been ongoing for several decades [1, 2, 3, 4, 5, 6]. Quark-gluon plasma (QGP) can be created in ultra-relativistic nucleus-nucleus interactions in Relativistic Heavy-Ion Collider (RHIC) experiments as well as in Large Hadron Collider (LHC) experiments. These experiments, notably the Beam Energy Scan (BES) program at RHIC in Brookhaven National Laboratory (BNL), offer insights into QGP properties and facilitate the mapping of the QCD phase diagram [6]. As one of the main goals of hot QCD physics [5, 6, 7, 8, 9, 10, 11], many efforts have been made for the searching for the CEP in the phase diagram with various methods and probes, such as the lattice calculations [12, 13, 14, 15], the ratio of viscosity to entropy density () [16, 17, 18, 19, 20], cumulants (skewness and kurtosis) [21, 22, 23, 24], conserved charge and baryon density fluctuations [25, 26, 28] as well as higher order moments [27, 29, 30, 31, 32], etc. In recent years, attention has also been payed to the ratio of light nuclei, which was proposed by Ref. [33, 34, 35] based on the coalescence model. Ones found that there exists a non-monotonic relation of the ratio at the STAR BES experiments [41], i.e. enhancements in the yield ratios relative to the coalescence baseline are observed in the most central collisions at 19.6 and 27 GeV, with a combined significance of 4.1. The enhancements are not observed in peripheral collisions or model calculations without critical fluctuation, and decreases with a smaller acceptance. The physics implications of these results on the QCD phase structure and the production mechanism of light nuclei in heavy-ion collisions are discussed [41]. As indicated by original references [33, 34], the non-monotonic relation could be related to the neutron density fluctuation as a function of center-of-mass energy [36, 37], and many interesting works on this direction have been demonstrated in the market [38, 37, 39, 40, 41, 42, 43]. In a recent STAR measurement of intermittency in Au + Au collisions at = 7.7-200 GeV at RHIC, the scaled factorial moments [44] of identified charged hadrons demonstrates that a non-monotonic energy dependence is observed in the 0-5% most central collisions with the scaling exponent reaching a minimum around = 27 GeV, which might be related to CEP [45]. In the study of heavy-ion collisions, hydrodynamics plays an important role [46, 47, 48]. Hydrodynamics, which is influenced by equations of state (EoS) as a function of energy and the baryon number densities, is applied to matter under local equilibrium during the process of heavy-ion collisions [49, 50, 51, 52]. However, the actual EoS of hot and dense QCD matter is still not so clear [46].
On the other hand, entropy is an important physical quantity to quantify how disorder of the system, and helpful to search the formation of the QGP state and investigate the correlations and event-by-event fluctuations [53, 54]. In the last century, Shannon introduced information entropy, extensively utilized across various disciplines. In nuclear physics, Ma pioneered the concept of multiplicity information entropy in 1999, applying it to heavy-ion collisions to study the phase transition of nuclear matter [56]. Subsequently, this concept has been widely explored in heavy-ion collision studies [57, 58]. Additionally, other forms of entropy, such as thermal entropy and entanglement entropy [58, 59, 60], have been employed to explore the formation of the QGP state. However, this study focuses only on multiplicity information entropy. Therefore, the objective of this paper is to extract information entropy from multiplicity distributions to investigate the effects of different Equations of State (EoS) in hydrodynamics and explore the phase transition of QGP matter in relativistic heavy-ion collisions.
II UrQMD model and information entropy
One of the purposes of the transport model for relativistic heavy-ion collisions is essentially for taking the effective solution of the relativistic Boltzmann equation [61, 46]
where is the distribution function of particle species and is the collision term. In the present work, simulations are based on the framework of the Ultra-relativistic Quantum Molecular Dynamics model (UrQMD), which includes about 60 baryonic species and 40 mesonic species, as well as their anti-particles [62, 63, 46, 64, 65]. The model includes particle rescattering, color string fragmentation, and the formation and decay of hadronic resonances. UrQMD is one of successful transport models for describing heavy-ion collisions across various energy ranges, spanning from the BNL Alternating Gradient Synchrotron (AGS) energies ( = 1 A10 A GeV) to SPS energies ( = 20 A160 A GeV), and extending to the full RHIC energy ( = 200 GeV) and the energy levels of the CERN Large Hadron Collider (up to 2.76 TeV for PbPb collisions) [66, 67, 68, 69, 65, 70, 71, 72, 73].
It should be noted that for the standard mode of UrQMD, the hadrons and strings can not be modeled at higher energy, such as 100 GeV. Thus, UrQMD combines a 3D1D ideal fluid dynamical simulation which is so-called a hybrid mode, including micro-Boltzmann transport and macro-hydrodynamics. For transforming the fluid dynamical fields to discrete hadrons, the Cooper-Frye equation,
is used. Here is the grand canonical Bose- or Fermi-distribution function, which depends on the local temperature and chemical potentials (x) [64]. In hydrodynamics, the EoS as a function of energy and the baryon number density is needed as an additional input to calculate the pressure, temperature, and chemical potential [46]. Three kinds of EoS, namely hadron gas (HG) EoS without deconfinement transition (EoS:HG) [74], chiral+hadron gas (CH) EoS with first order transition and critical endpoint (EoS:CH) [75], and bag model (BM) EoS with a strong first-order phase transition between QGP and hadronic phase (EoS:BM) [76], are employed in our simulations [46, 78, 77]. As a comparison, except for the hybrid mode, the standard (or default) mode is considered here. In UrQMD with hydrodynamics, as the system evolves and cools down, a particlization procedure is needed, which is dealt with in Cooper-Frye formalism. In this work, the default scenario known as the gradual particlization scenario (GF) is applied to the particlization. For the check, the proton number per rapidity as a function of center of mass energy is given in Fig. 1. The model results with different modes of UrQMD are compared to the data [37]. As shown by magenta line from EoS mode with BM, the values are the closest to the STAR data [37], and while red line which is the result of the default mode is also close to the STAR data. But the result with CH EoS is higher than the data when is above 20 GeV, and the HG EoS trend diverges beyond 10 GeV.
With the help of UrQMD, one can extract the information entropy in the relativistic heavy-ion collisions. The concept of multiplicity information entropy was initially introduced by Ma in Ref. [56], defined as:
where is the total number of a specific type of particle (e.g. hadrons or baryons etc.) produced in a simulation event and = is the normalized multiplicity probability. Here and denote event number of the multiplicity ‘’ and total event number, respectively. Here the bin width is set to ‘1’. It should be mentioned that = 1. Mathematically this entropy is called as the Shannon entropy [55]. In the present work, we extract information entropy in central 197Au+197Au collisions at impact parameter 03 fm with transverse momentum cut GeV/c and rapidity cut Y0.5.
III Results and discussion
Figure 2 illustrates the normalized multiplicity distribution of (a) hadrons, (b) anti-hadrons, (c) net-hadrons, (d) baryons, (e) anti-baryons, and (f) net-baryons at different center-of-mass energies in a framework of standard mode of UrQMD. It should be noted here that the hadrons and baryons denote positive particles. One can see that production of hadrons (or anti-hadrons or net-hadrons) increase with collision energy and the distributions become wider, but they are opposite for the baryons and net-baryons. And the multiplicity distribution of anti-baryons has the same behavior as the hadron one. One can easily obtain information entropy from these multiplicity distributions in Fig. 2. As shown in Fig. 3, we give the time evolution of information entropies of different particles. All information entropies of hadrons, anti-hadrons, net-hadrons, baryons, anti-baryons, and net-baryons increase with time and reach more than 90% of the total entropies within 05 fm/c. It indicates that the system reaches the chemical freeze-out fast within 05 fm/c. And compared to the information entropy of anti-hadrons, the hadron ones are higher. The same behavior exists between baryons and anti-baryons. This could be due to greater disorder or fluctuation for positive baryons and hadrons, since they are rich produced and then certainly induce greater fluctuation.
One can obtain information entropy at the final stage and give the relation between information entropy and center of mass energy. Although information entropy is calculated from the multiplicity probability as in Eq. (3), it is observed that information entropy is closely linked to the width of the multiplicity probability distribution. For instance, Fig. 4 illustrates the relation between information entropy and the log of full width at half maximum (FWHM) of with the standard UrQMD mode, revealing a linear relationship. This suggests that entropy really reflects the fluctuation or disorder within the multiplicity distribution; specifically, a wider multiplicity probability distribution corresponds to higher information entropy. We should note that here the width represents the multiplicity broadness. As the width increases, the shape of the multiplicity distribution becomes lower and lower, which within a certain range is closer to a uniform distribution, i.e., the more consistent the distribution of probabilities () of occurrence of the data. In an extreme case when the maximum entropy is reached, all corresponding probabilities of certain multiplicities tend to the same. In viewpoint of fluctuation, there is no uncertainity at , i.e. all have the same .
In our simulations, we explore four scenarios: without hydrodynamics (standard mode) and with hydrodynamic modes utilizing three different Equations of State (EoS): EoS:HG, EoS:CH, and EoS:BM, as shown in Fig. 5. In both of hadrons and anti-hadrons, information entropies increase with collision energy. It stems from that both the FWHMs of multiplicity distributions of hadrons and anti-hadrons increase as collision energy, which illustrates that increasing uncertainty for predicting how many hadrons’ multiplicity in a certain collision event.
In Fig. 5 (a), a comparison among those four cases is present, revealing that the information entropy is lowest for the scenario without hydrodynamics. It means that the FWHM for the w.o. hydro one is smaller and indicates that the hydro modes make the fluctuation larger since that for a distribution, the wider FWHM means larger standard deviation (or fluctuation). And the hydro mode with the EoS:BM is the largest one. However, the information entropy of the anti-hadrons with EoS:BM becomes close to the other two EoSs. If we compare the sensitivity of the information entropy (Fig. 5) or proton number (Fig. 1) to the UrQMD parameter set, we can say the information entropy is very sensitive to the hydo potential or not. For more details, Fig. 1 shows a very close proton number for the cases EoS:BM and w.o. hydro, but Fig. 5 shows a large difference of information entropy for the same cases.
Further, we give information entropies of baryons and anti-baryons, as shown in Fig. 6. One should notice that the behavior of baryons is different from the anti-baryons. For the latter with increasing collision energy, the information entropy increases monotonically as hadrons and anti-hadrons do. The information entropies of baryons, however, for all hydro cases with EoS exhibit valley shapes with a minimum at around 20 GeV, and consistent with a previous study utilizing a multiphase transport (AMPT) model in Ref. [82]. It indicates that the introduction of hydro mode induces additional fluctuation starting from 20 GeV. Furthermore, a comparison with the EoS:HG case, one can see that there is an enhancement around = 30 GeV for the information entropy for both EoS:CH and EoS:BM, which include first-order phase transitions as mentioned above. As one expects in relativistic heavy-ion collisions, large fluctuations of the baryon density would create at energies where the system passes through the first-order phase boundary or approaches to the CEP [6]. From the viewpoint of information entropy, large fluctuations of the baryon density could lead to broader multiplicity probability distribution giving larger information entropy. In other words, the enhancement at around 30 GeV collision energy, this non-monotonic behavior could correspond to the first-order phase boundary or CEP. For the anti-baryons, however, their information entropies only display monotonic increases as collision energy, as seen in Fig. 6 (b).
Now let us move on information entropies of net-baryons or net-protons. In previous works, fluctuations of the net-baryon number and net-proton number have been extensively discussed for the CEP in the viewpoints of higher order moments [79, 80, 5]. From the STAR measurement, the , where and represent kurtosis and variance respectively, of net-protons shows a nonmonotonic energy dependence, i.e. there appears a minimum which is close to zero around 20 GeV at 0-5 of most central Au+Au collisions [81]. The non-monotonic versus collision energy is qualitatively consistent with expectations from the QCD-based model expectation that, the higher the order of the moments is, the more sensitive it is to physics processes such as a critical point [21], or maybe an indication of onset the hydrodynamics process from the present work. In addition, as mentioned before, enhancements of the ratio relative to the coalescence baseline as well as a minimum scale exponent of intermittency at the STAR BES experiments are also observed in the most central collisions at around 27 GeV [41]. Here the information entropies for multiplicity distributions of net-baryons and net-protons are investigated and displayed in Fig. 7. From Fig. 7(a), one can see that information entropy with EoS:HG shows smooth line as the result of baryons. However, both information entropies of EoS:BM and EoS:CH show enhancements around 30 GeV. In addition, it can be noted that the enhanced region with EoS:BM is broader than that from EoS:CH which could be due to a strong first-order phase transition in the EoS:BM hydro mode.
IV Conclusions
In summary, we simulated central 197Au+197Au collisions in the framework of UrQMD model which considers a standard mode or hybrid modes with three kinds of EoSs, i.e. EoS:HG without deconfinement transition, EoS:CH with the first-order phase transition and critical endpoint, and EoS:BM with a strong first-order phase transition between QGP and hadronic phase. By calculating the multiplicity information entropies of hadrons, anti-hadrons, baryons, anti-barons, net-baryons and net-protons, we explore the possible phase transition and EoS effects for the hydrodynamics. The dependence of information entropy on collision energy reflects the relationship between the width of multiplicity probability distribution of particles and collision energy. With the help of information entropies of baryons and net-baryons, it is found that both information entropies have valleys at around = 20 GeV with hydro hybrid modes, which indicates the onset of hydrodynamics behavior in the model. In addition, an enhancement emerges around = 30 GeV with EoS:CH and EoS:BM which is consistent with recent experimental observations from as well as intermittency analysis which give enhancement around =27 GeV, that may suggest the CEP. In one word, the present work suggests information entropy could be taken as an alternative observable on the QGP phase transition, and we expect that it can be checked from the experimental side.
Acknowledgements.
Authors thank Dr. Chen Zhong for maintaining the high-quality performance of Fudan supercomputing platform for nuclear physics and the discussions with Dr. Kai-Jia Sun. This work received partial supports from the National Natural Science Foundation of China under Contract Nos. , , , ,and , the Strategic Priority Research Program of CAS under Grant No. XDB34000000, the Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008,the STCSM under Grant No. 23590780100, and the Natural Science Foundation of Shanghai under Grant No. 23JC1400200.References
- [1] N. Cabibbo and G. Parisi, Phys. Lett. B 59, 67 (1975).
- [2] G. Baym, Nucl. Phys. A 698, xxiii-xxxii (2002).
- [3] K. Fukushima, J. Phys. G: Nucl. Part. Phys. 35, 104020 (2008).
- [4] K. Fukushima and T. Hatsuda, Rep. Prog. Phys. 74, 014001 (2011).
- [5] X. F. Luo and N. Xu, Nucl. Sci. Tech. 28, 112 (2017).
- [6] A. Bzdak et al., Phys. Rep. 853, 1–87 (2020).
- [7] Y. L. Du, C. M. Li, C. Shi, et al., Nucl. Tech. (in Chinese) 46, 040009 (2023).
- [8] F. P. Li, L. G. Pand, X. N. Wang, Nucl. Tech. (in Chinese) 46, 040014 (2023) .
- [9] Wan-Bing He, Yu-Gang Ma, Long-Gang Pang, Hui-Chao Song, Kai Zhou, Nucl. Sci. Tech. 34, 88 (2023).
- [10] Yu-Gang Ma, Long-Gang Pang, Rui Wang, Kai Zhou, Chin. Phys. Lett. 40, 122101 (2023) .
- [11] Qian Chen, Guo-Liang Ma and Jin-Hui Chen, Nucl. Tech. (in Chinese), 46, 040013 (2023).
- [12] Z. Fodor and S. D. Katz, JHEP 03, 014 (2002).
- [13] P. de Forcrand and O. Philipsen, Nucl. Phys. B 673, 170 (2003).
- [14] H. T. Ding, S.T. Li, J.H. Liu, Nucl. Tech. (in Chinese) 46, 040008 (2023).
- [15] Z. R. Zhu, Y. Q. Zhao, D. F. Hou, Nucl. Tech. (in Chinese) 46, 040007 (2023).
- [16] D. Teaney, Phys. Rev. C 68, 034913 (2003).
- [17] S. Gavin and M. Abdel-Aziz, Phys. Rev. Lett. 97, 162302 (2006).
- [18] J. E. Bernhard, J. S. Moreland, S. A. Bass, Nat. Phys. 15, 1113-1117 (2019).
- [19] X. G. Deng, P. Danielewicz, Y. G. Ma, H. Lin, Y. X. Zhang, Phys. Rev. C 105, 064613 (2022).
- [20] X. G. Deng, D. Q. Fang, Y. G. Ma, Prog. Part. Nucl. Phys., 136, 104095 (2024).
- [21] M. A. Stephanov, Phys. Rev. Lett. 102, 032301 (2009).
- [22] M. A. Stephanov, Phys. Rev. Lett. 107, 052301 (2011).
- [23] C. Herold, M. Bleicher, M. Nahrgang et al., Eur. Phys. J. A 54, 19 (2018).
- [24] J. M. Torres-Rincon and E. Shuryak, Proceedings of Science 347, 176 (2019).
- [25] C. Sasaki, B. Friman, and K. Redlich, Phys. Rev. Lett. 99, 232301 (2007).
- [26] M. A. Stephanov, Phys. Rev. Lett. 102, 032301 (2009).
- [27] M. A. Stephanov, S. Ejiri, and M. Kitazawa, Phys. Rev. Lett. 103, 262301(2009).
- [28] C. Liu, X.G. Deng, Y.G. Ma, Nucl. Sci. Techn. 33, 52 (2022).
- [29] J. Steinheimer and J. Randrup, Phys. Rev. Lett. 109, 212301 (2012).
- [30] C. M. Ko and F. Li, Nucl. Sci. Tech. 27, 140 (2016).
- [31] F. Li and C. M. Ko, Phys. Rev. C 95, 055203 (2017).
- [32] Ru-Xin Cao, Song Zhang, Yu-Gang Ma, Phys. Rev. C 106, 014910 (2022).
- [33] K. J. Sun, L. W. Chen, C. M. Ko, Z. Xu, Phys. Lett. B 774, 103 (2017).
- [34] K. J. Sun, L. W. Chen, C. M. Ko, J. Pu, Z. Xu, Phys. Lett. B 781, 499 (2018).
- [35] K. J. Sun, L. W. Chen, C. M. Ko, et al., Nucl. Tech. (in Chinese) 46, 040012 (2023).
- [36] Dingwei Zhang (STAR Collaboration), Nucl. Phys. A 1005, 121825 (2021).
- [37] Hui Liu, Dingwei Zhang, Shu He, Kai-jia Sun, Ning Yu, Xiaofeng Luo, Phys. Lett. B 805, 135452 (2020).
- [38] E. Shuryak and J. M. Torres-Rincon, Phys. Rev. C 101, 034914 (2020).
- [39] K. J. Sun and C. M. Ko, arXiv:2005.00182 [nucl-th] (2020).
- [40] X. G. Deng and Y. G. Ma, Phys. Lett. B 808, 135668 (2020).
- [41] M. I. Abdulhamid et al. (STAR Collaboration), Phys. Rev. Lett. 130, 202301 (2023).
- [42] C. M. Ko, Nucl. Sci. Tech. 34, 80 (2023).
- [43] Kai-Jia Sun, Rui Wang, Che Ming Ko, Yu-Gang Ma, Chun Shen, Nature Commun. 15, 1074 (2024)
- [44] Y. G. Ma, Eur. Phys. J. A 30, 227-242 (2006).
- [45] M. I. Abdulhamid et al. (The STAR Collaboration), Phys. Lett. B 845, 138165 (2023).
- [46] H. Petersen, J. Steinheimer, G. Burau, M. Bleicher, H. Stöcker, Phys. Rev. C 78, 044901 (2008).
- [47] J. Steinheimer et al., Phys. Rev. C 77, 034901 (2008).
- [48] S. Gope and B. Bhattacharjee, Eur. Phys. J. A 57, 44 (2021).
- [49] H. C. Song, Y. Zhou, K. Gajdoov, Nucl. Sci. Tech. 28, 99 (2017).
- [50] Chun Shen, Li Yan, Nucl. Sci. Tech. 31, 112 (2020).
- [51] Z. Yang, T. Luo, W. Chen et al., Phys. Rev. Lett. 130, 052301 (2023).
- [52] Guang-You Qin, Nucl. Sci. Tech. 34, 22, (2023).
- [53] K. Yamazaki, T. Matsui, and G. Baym, Nucl. Phys. A 933, 245255, (2015).
- [54] C. Herold et al., Phys. Lett. B 790, 557–562 (2019).
- [55] C. E. Shannon, The Mathematical Theory of Communication, The University of Illinois Press, Urbana, Ill, USA, 1948.
- [56] Y. G. Ma, Phys. Rev. Lett. 83, 3617 (1999).
- [57] Chun-Wang Ma, Yu-Gang Ma, Prog. Part. Nucl. Phys. 99, 120 (2018).
- [58] F. Li and G. Chen, Eur. Phys. J. A 56:167 (2020).
- [59] M. Reiter et al, Nucl. Phys. A, 643, 29112, (1998).
- [60] A. Kovner, M. Lublinsky, and M. Serino, Phys. Lett. B 792, 4–15 (2019).
- [61] S. R. De Groot, W. A. Van Leeuwen, and C. G. Van Weert, Relativistic Kinetic theory: Principles and Applications (North-Holland, Amsterdam, 1980).
- [62] S. A. Bass, M. Belkacem, M. Bleicher et al., Prog. Part. Nucl. Phys. 41, 225 (1998).
- [63] M. Bleicher, E. Zabrodin, C. Spieles et al., J. Phys. G 25, 1859 (1999).
- [64] J. Steinheimer, V. Vovchenko, J. Aichelin et al., EPJ Web of Conferences 171, 05003 (2018).
- [65] F. H. Qiao, X. G. Deng, Y. G. Ma, Phys. Lett. B 850, 138535 (2024).
- [66] M. Mitrovski, T. Schuster, G. Gräf et al., Phys. Rev. C 79, 044901 (2009).
- [67] P. P. Bhaduri and S. Chattopadhyay, Phys. Rev. C 81, 034906 (2010) .
- [68] S. Sombun, K. Tomuang, A. Limphirat et al., Phys. Rev. C 99, 014901 (2019).
- [69] Zepeng Gao, Qingfeng Li, Nucl. Tech. 46, 080009 (2023).
- [70] Bo Gao, Yongjia Wang, Qingfeng Li, Baochun Li, Nucl. Tech. 46, 070501 (2023).
- [71] Kui Xiao, Pengcheng Li, Yongjia Wang, Fuhu Liu, Qingfeng Li, Nucl. Sci. Tech. 34, 62 (2023).
- [72] S. W. Lan, S. S. Shi, Nucl. Sci. Tech. 33, 21 (2022).
- [73] P. C. Li, Y. J. Wang, Q. F. Li, H. F. Zhang, Sci. Chin. Phys. Mech. and Astro. 66, 222011 (2023).
- [74] D. Zschiesche, S. Schramm, J. Schaffner-Bielich, et al., Phys. Lett. B 547, 7 (2002).
- [75] J. Steinheimer, S. Schramm and H. Stöcker, Phys. Rev. C 84, 045208 (2011)
- [76] D. H. Rischke, Y. Pürsün and J. A. Maruhn, Nucl. Phys. A 595, 383 (1995).
- [77] Yu. B. Ivanov and A. A. Soldatov, Eur. Phys. J. A 52: 367 (2016).
- [78] Q. F. Li, J. Steinheimer, H. Petersen, et al., Phys. Lett. B 674, 111-116 (2009).
- [79] K. Morita, B. Friman, K. Redlich, Phys. Lett. B 714,178–183 (2015).
- [80] C. Herold et al., Phys. Rev. C 93, 021902(R) (2016).
- [81] M. S. Abdallah et al. (STAR Collaboration), Phys. Rev. C 104, 024902 (2021).
- [82] J. Pu, Y. B,Yu, K. X. Kai, et al., Phys. Lett. B 841, 137909 (2023).